Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y = mx + b
(-10,8); 7x-5y = 2
I need step by step instructions on solving this. I have worked it several times but when I get to the part of removing the parentheses and simplifying, I get lost.
Please help. Thank you!
Tracy
email: tgraham@fkp.com

Answer by stanbon(75887)

Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y = mx + b
(-10,8); 7x-5y = 2
----
The given line has slope = 7/5
---
Find y = mx +b where m = 7/5 and y = 8 when x = -10:
8 = (7/5)(-10) + b
8 = -35 + b
b = 43
===============
Equation:
y = (7/5)x + 43
===================
Cheers,
Stan H.